OpenStudy (anonymous):

Bill is trying to figure out the area of his square yard. He knows that one side of the yard is equal to the length of his garage. The dimensions of his rectangular garage are 19 by ten. What is the area of the yard?

OpenStudy (mathstudent55):

Since the yard is a square, both the length and the width are the same. The side of the square yard is the same as the length of the garage. The garage has a length of 19 and a width of 10. That means the yard has a side of 19. How do you find the area of a square?

OpenStudy (anonymous):

Area=w*h

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Right. For a square, w = h = s A = s * s The square yard has a side of 19. A = s * s = 19 * 19

OpenStudy (anonymous):

361?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

@mathstudent55 Thank you!

OpenStudy (mathstudent55):

wlcm